OpenStudy (anonymous):

MY ANSWER IS D. PLEASE CHECK TO SEE IF I'M RIGHT? VERY IMPORTANT!!!!!!! If you were to use the elimination method to solve the following system, choose the new system of equations that would result after the variable y is eliminated in the first and second equations, then the second and third equations. 3x - y + 3z = 1 4x + y - 6z = 3 4x - 2y + 3z = -3 (4 points) Answer 7x - 3z = 4 12x - 9z = 3 7x - 3z = 4 4x - 6z = 3 3x + 3z = 1 12x - 9z = 3 3x + 3z = 1 4x - 6z = 3

OpenStudy (anonymous):

Hey. add eqs 1 & 2 then u get 7x - 3z = 4 multiply eq-2 by 2 and then add to eq-3, as below 2*(4x + y - 6z) = 2*3 => 8x + 2y - 12z = 6, now add to eq-3, then u get 12x - 9z = 3 so, the correct answer is A option..

OpenStudy (anonymous):

Did, this help???